Student population of a university has 30% Asian, 40% American, 20% European

and 10% African students. It is known that 40% of all Asian students, 50% of all

American students, 60% of all European students and 20% of all African students are

girls. Find the probability that a girl chosen at random from the university is an Asian.


1
Expert's answer
2021-09-12T23:52:23-0400

Let "A_1" denote asian student, "A_2" denote american student, "A_3" denote european student, "A_4" denote african student, and "G" denote that a student is a girl.

Given

"P(A_1)=0.3,"

"P(A_2)=0.4,"

"P(A_3)=0.2,"

"P(A_4)=0.1,"

"P(G|A_1)=0.4,"

"P(G|A_2)=0.5,"

"P(G|A_3)=0.6,"

"P(G|A_4)=0.2"

By the Law of the Total Probability


"P(G)=P(A_1)P(G|A_1)+P(A_2)P(G|A_2)+""+P(A_3)P(G|A_3)+P(A_4)P(G|A_4)=""=0.3(0.4)+0.4(0.5)+0.2(0.6)+0.1(0.2)=0.46"

Use the Bayes' Theorem


"P(A_1|G)=\dfrac{P(A_1)P(G|A_1)}{P(G)}=""=\dfrac{0.3(0.4)}{0.46}={6\over 23}\approx0.260870"